Find the value of
54
Find the value of
54
Find the value of
54

Grade:12th pass

2 Answers

Harshit Singh
2 years ago
Dear Student

The above value can be written as,
= 5 × 5 × 5 × 5
= 625

Thanks
Shraddha Singh
2 years ago
Dear Student,
 
The value of the following number can be calculated as
5⁴ = 5×5×5×5 = 625
 
Hope it helps!
Thank you ✌️